Rios Financial Co. is a regional insurance company that began operations on January 1, Year 1. The following transactions relate to trading securities acquired by Rios Financial Co., which has a fiscal year ending on December 31:

Year 1

Feb. 1. Purchased 4,900 shares of Caldwell Inc. as a trading security at $35 per share plus a brokerage commission of $490.
May 1. Purchased 1,800 shares of Holland Inc. as a trading security at $47 plus a brokerage commission of $198.
July 1. Sold 2,550 shares of Caldwell Inc. for $33 per share less a $95 brokerage commission.
31. Received an annual dividend of $0.30 per share on Caldwell Inc. stock.
You are in Column Year 1Dec. 31. The portfolio of trading securities was adjusted to fair values of $33 and $46 per share for Caldwell Inc. and Holland Inc., respectively.

Year 2

Apr. 1. Purchased 4,800 shares of Fuller Inc. as a trading security at $26 per share plus a $240 brokerage commission.
July 31. Received an annual dividend of $0.40 per share on Caldwell Inc. stock.
Oct. 14. Sold 960 shares of Fuller Inc. for $28 per share less a $30 brokerage commission.
You are in Column Year 2Dec. 31 The portfolio of trading securities had a cost of $267,315 and a fair value of $346,935, requiring a debit balance in Valuation Allowance for Trading Investments of $79,620 ($346,935 - $267,315). Thus, the credit balance from December 31, Year 1, is to be adjusted to the new balance.
Required:
1. Journalize the entries to record these transactions. Round all final amounts to the nearest whole dollar.*
2. Prepare the investment-related current asset balance sheet presentation for Rios Financial Co. on December 31, Year 2.*
3. How are unrealized gains or losses on trading investments presented in the financial statements of Rios Financial Co.?
*Refer to the information given and the Chart of Accounts and Amount Descriptions provided for the exact wording of the answer choices for text entries. “Less” or “Plus” will automatically appear if it is required. For those boxes in which you must enter subtracted or negative numbers use a minus sign.

none

X

Chart of Accounts

CHART OF ACCOUNTS
Rios Financial Co.
General Ledger
ASSETS
110 Cash
111 Petty Cash
120 Accounts Receivable
121 Allowance for Doubtful Accounts
131 Notes Receivable
132 Interest Receivable
141 Merchandise Inventory
145 Office Supplies
146 Store Supplies
151 Prepaid Insurance
161 Investments-Caldwell Inc.
162 Investments-Holland Inc.
163 Investments-Fuller Inc.
165 Valuation Allowance for Trading Investments
166 Valuation Allowance for Available-for-Sale Investments
181 Land
191 Store Equipment
192 Accumulated Depreciation-Store Equipment
193 Office Equipment
194 Accumulated Depreciation-Office Equipment
LIABILITIES
210 Accounts Payable
221 Notes Payable
231 Interest Payable
241 Salaries Payable
251 Sales Tax Payable
EQUITY
311 Common Stock
312 Paid-In Capital in Excess of Par-Common Stock
321 Preferred Stock
322 Paid-In Capital in Excess of Par-Preferred Stock
331 Treasury Stock
332 Paid-In Capital from Sale of Treasury Stock
340 Retained Earnings
350 Unrealized Gain (Loss) on Available-for-Sale Investments
351 Cash Dividends
352 Stock Dividends
390 Income Summary
REVENUE
410 Sales
611 Interest Revenue
612 Dividend Revenue
631 Gain on Sale of Investments
641 Unrealized Gain on Trading Investments
EXPENSES
511 Cost of Merchandise Sold
512 Bad Debt Expense
515 Credit Card Expense
516 Cash Short and Over
520 Salaries Expense
531 Advertising Expense
532 Delivery Expense
533 Repairs Expense
534 Selling Expenses
535 Rent Expense
536 Insurance Expense
537 Office Supplies Expense
538 Store Supplies Expense
561 Depreciation Expense-Store Equipment
562 Depreciation Expense-Office Equipment
590 Miscellaneous Expense
710 Interest Expense
731 Loss on Sale of Investments
741 Unrealized Loss on Trading Investments

none

X

Amount Descriptions

Amount Descriptions

In CengageNOW, an Amount Description is a text entry other than an Account that has an amount associated with it.

Increase in retained earnings
Net income
Net loss
Other comprehensive income (loss)
Other income (loss)
Retained earnings, December 31, Year 2
Retained earnings, January 1, Year 2
Trading investments (at cost)
Trading investments (at fair value)
Answer #1

Journal entries for purchase of 4900 shares $35 per share plus a brokerage commission of $490 In Company C:

= (4900*35) + 490 = 171990

Hence,

Investments in Company C 171990 Dr.

To Cash 171990 Cr.

Journal entries for purchase of 1800 shares $47 per share plus a brokerage commission of $198 In Company H:

= (1800*47) + 198 = 84798

Hence,

Investments in Company C 84798 Dr.

To Cash 84798 Cr.

Journal entries for purchase of 2550 shares $33 per share plus a brokerage commission of $95 In Company C:

= (2550*33) - 95 = 84055

Cost of stock investment sold = 2550 shares * { 171990/4900} = 89505 which is investment

July 1 Cash 84055

Loss on sale of investment 5450

To Investment 89505

Journal entries for Dividend issued in Company C

Dividend received = Number of shares * dividend per share.

= 4900-4500 * 0.3

  = 120

Cash Dr. 120

To Dividend revenue Cr. 120

Journal entries for Fuller Company for Year 2 :

= (4800 shares * 26 ) + 240

= 125040

Investment in Fuller 125040 Dr.

To Cash 125040 Cr.

Journal entries for Dividend issued Company C:

Dividend received = Number of shares * dividend per share.

= 4900-4500 * 0.4

  = 160

Cash Dr. 160

To Dividend revenue Cr. 160

Sold 960 shares of Fuller Inc. for $28 per share less a $30 brokerage commission.

Cash ( 960*28) - 30 26850

Loss on sale of investment 6846

To investments 35.1 * 960 33696

PARTIAL BALANCE SHEET:

Trading investment 267315

Add valuation allowance 79620

Trading investment at Fair value 346935

Part 3:

At the end of year 1, the unrealized losses and gains on trading investments will be reported as other losses at $5450.